OpenStudy (anonymous):

complete the square. x^2-12x=-28 explain...

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

You take the -12 and get half of it. half of -12 is -6 Then you square that. -6 squared is 36, the negatives cancel. Add that to both sides. now you have x^2 -12x + 36 = -28 + 36! Factor. (x+6)^2 = 8 So then x + 6 is the square root of 8. simplify, and then you get